The following is a partial performance report for a revenue center for the Northern Division of Summer Food . Restaurants.

Summer Food Restaurants

Sales Revenue - Northern Division

For the Month Ended June 30

Variance

Product ActualSales Budgeted Sales Variance Variance Percentage

Food $159,650 $155,000 . ? ?

Dessert 21,340 22,000 . ? ?

Bar 59,160 . 58,000 ? ?

Catering 57,000 60,000 ? ?

Fill in the missing amounts. Indicate whether each variance is favorable​ (F) or unfavorable​ (U). ​(Enter the variances as positive numbers and round the variance percentages to the nearest tenth of a​ percent, X.X%. Label each variance as favorable​ (F) or unfavorable​ (U).)

Summer Food Restaurants

Northern Division - Sales Revenue for

For the Month Ending June 30

Variance

Product

Actual Sales

Budgeted Sales

(U or F)

Food

$159,650

$155,000

Dessert

$21,340

$22,000

Bar

$59,160

$58,000

Catering

$57,000

$60,000

Choose from any list or enter any number in the input fields and then click Check Answer.
Answer #1

Summer Food Restaurants

Northern Division - Sales Revenue for

For the Month Ending June 30

Variance

Product

Actual Sales

Budgeted Sales

(U or F)

Variance %

Food

$159,650

$155,000

4,650 F 4,650/155,000 = 3%

Dessert

$21,340

$22,000

660 U 660/22,000 = 3%

Bar

$59,160

$58,000

1,160 F 1,160/58,000 = 2%

Catering

$57,000

$60,000

3,000 U 3,000/60,000 = 5%

Kindly comment if you need further assistance. Thanks